# Perfect Matching in Product Graphs and in their Random Subgraphs

### v2

@@ 3.3 Typically no obstructions
using $\frac{2C^{2}}{\ln C}\geq 11$ for $C\geq 2$ and $\frac{2\ln n}{2^{d/(100C)^{3}}}=o(\ln d)$. A union bound over the less than $n$ possible values of $u$ completes the proof. ∎
We now turn to consider obstructions with large $U$. We consider several cases separately. The proofs in Lemmas 3.9-3.14 follow quite similar calculations, however, each such calculation requires a slightly different approach. As these differences are telling of the different challenges in showing that such typical obstructions do not appear, we left the calculations explicit.
Obstructions with $s=b=0$.
@@
Whp, there are no obstructions with $s=b=0$.
@@
Fix $u$ and $w$, and let $\mathcal{B}_{u,w}$ be the event that there is a minimal obstruction with $u=|U|,w=|W|$ and $s=b=0$. By Lemma 3.3, there are at most $2n^{w/2}(ed)^{w/2}$ such minimal obstructions. In particular, $\ell=|V_{1}|=n-u-w$.
We have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq d\ell+w(d-1).$
As there are at most $du$ edges touching $U$ in $G$, we have by the union bound
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}u\in[1,n/2-1]\\ w\in[0,n-2u-1]\end{subarray}}\mathcal{B}_{u,w}\right)$ $\leq\sum_{u=1}^{n/2-1}\sum_{w=0}^{n-2u-1}2n^{w/2}(ed)^{w/2}(1-p)^{d(\ell+w-u)-Cw}$
$\leq\sum_{u=1}^{n/2-1}\sum_{w=0}^{n-3}2n^{-(1-\epsilon)(\ell-u)-w/3}.$
Recall that since $u$ and $w$ are fixed, $\ell-u$ is fixed as well and ranges from $2$ to $n-2$. Furthermore, note that the sum over $w$ is a geometric sum and is at most twice its value when $w=0$. Thus, we have that
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}u\in[1,n/2-1]\\ w\in[0,n-2u-1]\end{subarray}}\mathcal{B}_{u,w}\right)$ $\leq 4\sum_{j=2}^{n-2}n^{-(1-\epsilon)j}=o(1),$
as required. ∎
Obstructions with $s\neq 0,b=0$.
Lemma 3.10.
Whp, there are no obstructions with $s\neq 0$ and $b=0$.
Proof.
Fix $u,w$ and $s$, and let $\mathcal{B}_{u,w,s}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $0\neq s=|S|$ and $b=0$. We then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{j=w/2}e(W_{j},W_{j}^{C})\geq d\ell_{1}+ds-2Cs\log_{2}d+w(d-1),$
since the components in $S$ are of size at most $d^{2}$, and thus by Theorem 1.2, any $K_{i}$ of size at most $d^{2}$ has that $e(K_{i},K_{i}^{C})\geq|K_{i}|(d-2C\log_{2}d)$.
By Lemma 3.3 there are at most $2n^{w/2+\ell_{2}}(ed)^{w/2+2s}$ choices for such a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $b=0$. As before, we obtain that
$\mathbb{P}\left(\bigcup_{w\in[0,n-u-s-\ell_{1}]}\mathcal{B}_{u,w,s}\right)$ $\leq\sum_{w=0}^{n-u-s-\ell_{1}}2n^{w/2+\ell_{2}}(ed)^{w/2+2s}(1-p)^{d(\ell_{1}+s+w-u)-2Cs\log_{2}d-w}$
$\leq 4n^{\ell_{2}}(ed)^{2s}(1-p)^{d(\ell_{1}+s-u)-2Cs\log_{2}d}.$
Now, $\ell_{1}+\ell_{2}\geq u+1$, and thus $\ell_{1}-u\geq 1-\ell_{2}$. Hence,
$n^{\ell_{2}}(ed)^{2s}(1-p$ $)^{d(\ell_{1}+s-u)-2C\log_{2}ds}\leq n^{\ell_{2}}(ed)^{2s}(1-p)^{d(s-\ell_{2}+1)-2Cs\log_{2}d}$
$\leq n^{-(1-\epsilon)(s-2\ell_{2}+1)}(ed)^{2s}(1-p)^{-2Cs\log_{2}d}$
$\leq\exp\left\{-(1-\epsilon)(s-2\ell_{2}+1)\ln n+5C\left(1+\ln\left(\frac{1}{1-p}\right)\right)s\ln d\right\}.$
Note that the above decreases as $s$ increases, and therefore, using $s\geq 3\ell_{2}$ and $\ell_{2}\geq\frac{s}{d^{2}}$, we have that
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}s\in[3,n-u-\ell_{1}]\\ w\in[0,n-u-s-\ell_{1}]\end{subarray}}\mathcal{B}_{u,w,s}\right)$ $\leq 4\sum_{s=3}^{n-u-\ell_{1}}\exp\left\{-(1-\epsilon)(\ell_{2}+1)\ln n+15C\left(1+\ln\left(\frac{1}{1-p}\right)\right)\ell_{2}\ln d\right\}$
$\leq 4d^{2}\sum_{\ell_{2}=1}^{n}\exp\left\{-(1-2\epsilon)(\ell_{2}+1)\ln n\right\}$
$\leq 5d^{2}\exp\left\{-(1-2\epsilon)2\ln n\right\}=o(1/n),$
where we used the fact that $p$ is bounded away from one. Therefore, by the union bound over the at most $n$ values of $u$, we have that whp there is no obstruction violating the statement of the Lemma. ∎
Obstructions with $0\neq b\leq\frac{n}{2}$.
We first treat this case under the assumption that the edge boundary of $B$ in $G$ is not too small.
Lemma 3.11.
Whp there are no obstructions with $0\neq b$ and $e(B,B^{C})\geq b\ln^{2}d$.
Proof.
Fix $u,w,s,\ell_{2},b$, and $\ell_{3}$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $0\neq b=|B|$. Similarly to before, we then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w.$
Since we assume that $e(B,B^{C})\geq b\ln^{2}d$, we obtain that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+b\ln^{2}d+wd-w.$
Now, by Lemma 3.3, there are at most $2n^{w/2+\ell_{2}+2\ell_{3}}(ed)^{w/2+2s+b}=2n^{\ell-\ell_{1}+\ell_{3}+w/2}(ed)^{2s+b+w/2}$ such minimal obstructions with $u=|U|,w=|W|,s=|S|$ and $b=|B|$. As before, we obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2n^{\ell-\ell_{1}+\ell_{3}}(ed)^{2s+b}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+b\ln^{2}d}\cdot n^{w/2}(ed)^{w/2}(1-p)^{dw-w}$
$\leq 2n^{\ell-\ell_{1}+\ell_{3}}(ed)^{2s+b}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+b\ln^{2}d}\cdot n^{-w/3}$
$\leq 2n^{\ell-\ell_{1}+\ell_{3}}(ed)^{2s+b}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+b\ln^{2}d}$
$\leq 2\exp\left\{(\ell-\ell_{1}+\ell_{3})\ln n+(2s+b)\ln(ed)-pb\ln^{2}d\right\}n^{-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)}$
$\leq 2\exp\left\{(\ell-\ell_{1})\ln n+2s(\ln d+1)-\frac{p}{2}b\ln^{2}d-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)\ln n\right\}$
$\leq 2\exp\left\{(\ell+\ell_{3}-(2-\epsilon)\ell_{1}+(1-\epsilon)u)\ln n-s(1-2\epsilon)\ln n-\frac{p}{2}b\ln^{2}d\right\}.$
Recall that $\ell=\ell_{1}+\ell_{2}+\ell_{3}$ and $\ell_{2}\leq\frac{s}{3}$ and $\ell_{3}\leq\frac{b}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{(\ell+\ell_{3}-(2-\epsilon)(\ell-\ell_{2}-\ell_{3})+(1-\epsilon)u)\ln n-s(1-2\epsilon)\ln n-\frac{p}{2}b\ln^{2}d\right\}$
$\leq$ $2\exp\left\{\left((1-\epsilon)\left(u-\ell\right)+\frac{2s}{3}+\frac{3b}{d^{2}}\right)\ln n-s(1-2\epsilon)\ln n-\frac{p}{2}b\ln^{2}d\right\}$
$\leq$ $2\exp\left\{(1-\epsilon)(u-\ell)\ln n+O\left(\frac{b}{d}\right)-s\left(1-2\epsilon-\frac{2}{3}\right)\ln n-\frac{p}{2}b\ln^{2}d\right\}$
$\leq$ $2\exp\left\{O\left(\frac{b}{d}\right)-\frac{p}{2}b\ln^{2}d\right\},$
@@
where in the last step we used that $u<\ell$. Since $b\geq d^{2}$, by the union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$, we have that the probability of an obstruction violating the statement of the claim is at most
@@
Whp there are no obstructions with $0\neq b\leq\frac{n}{2}$ and $e(B,B^{C})\leq b\ln^{2}d$.
@@
Note that every set $B\subseteq V$ with $e(B,B^{C})\leq b\ln^{2}d$ is of size at least $\frac{n}{d^{\ln^{2}d}}$ since by 1.2, for $b<\frac{n}{d^{\ln^{2}d}}$,
$e(B,B^{C})\geq\frac{b\log_{C}(n/b)}{C}\geq\frac{b\ln^{2}d\log_{C}d}{C}>b\ln^{2}d$
Fix $u,w,s,\ell_{2},b$, and $\ell_{3}$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $0\neq b=|B|\leq\frac{n}{2}$. Similarly to before, we then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+f(b)+wd-w.$
@@
By the proof of Lemma 3.3, there are at most $n^{\ell_{2}+w/2}(ed)^{2s+w/2}$ ways to choose $W\cup S$. Furthermore, since we assume $e(B,B^{C})\leq b\ln^{2}d$ and $b\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2b}{\ln d}\right\}$ choices for $B$. Therefore, by Lemma 3.2, there are at most $2n^{\ell_{2}+w/2}(ed)^{2s+w/2}\exp\left\{\frac{2b}{\ln d}\right\}$ such minimal obstructions. We obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2b}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(b)}\cdot n^{w/2}(ed)^{w/2}(1-p)^{dw-w}$
$\leq 2n^{\ell_{2}}(ed)^{2s}\exp\left\{\frac{2b}{\ln d}\right\}(1-p)^{d(\ell_{1}+(1+o(1))s-u)+f(b)}$
$\leq 2\exp\left\{\ell_{2}\ln n+2s\ln(ed)+\frac{2b}{\ln d}-pf(b)\right\}n^{-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)}$
$\leq 2\exp\left\{\ell_{2}\ln n+2s(\ln d+1)+\frac{2b}{\ln d}-pf(b)-(1-\epsilon)(\ell_{1}+(1+o(1))s-u)\ln n\right\}$
$\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell_{1}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2b}{\ln d}-pf(b)\right\}.$
Recall that $\ell=\ell_{1}+\ell_{2}+\ell_{3}$ and $\ell_{2}\leq\frac{s}{3}$ and $\ell_{3}\leq\frac{b}{d^{2}}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{(\ell_{2}-(1-\epsilon)(\ell-\ell_{2}-\ell_{3}-u))\ln n-s(1-2\epsilon)\ln n+\frac{2b}{\ln d}-pf(b)\right\}$
$\leq 2\exp\left\{\left(\epsilon\frac{s}{3}+(1-\epsilon)\frac{b}{d^{2}}-(1-\epsilon)(\ell-u)\right)\ln n-s(1-2\epsilon)\ln n+\frac{2b}{\ln d}-pf(b)\right\}$
$\leq 2\exp\left\{\frac{3b}{\ln d}-pf(b)\right\},$
@@
where in the last step we used that $u<\ell$. Since we assume that $b\leq\frac{n}{2}$ and $n\leq C^{d}$, we have that $\log_{C}b\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f(b)\geq\frac{b}{C^{2}}$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w,s,b}\right)$ $\leq 2\exp\left\{b\left(\frac{3}{\ln d}-\frac{p}{C^{2}}\right)\right\}\leq\exp\left\{-\frac{d^{2}}{2C^{2}}\right\}.$
union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},n$ and $\ell_{3}$ completes the proof. ∎
Obstructions with $b>\frac{n}{2}$.
@@
Note that here, by Lemma 3.8, we may assume that $u\geq\frac{n}{d^{C^{3}/p}}$, and thus $|B^{C}|\geq\frac{n}{d^{C^{3}/p}}$. We begin by assuming that $B^{C}$ has a large edge-boundary.
@@
Whp there are no obstructions with $b>\frac{n}{2}$ and $e(B^{C},B)\geq|B^{C}|\ln^{2}d$.
@@
Fix $u,\ell_{1},w,s$ and $b$. We then have that
$\sum_{i=1}^{\ell}e(K_{i},K_{i}^{C})+\sum_{j=1}^{w/2}e(W_{j},W_{j}^{C})\geq\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w.$
By Claim 3.6, we may assume that $\ell_{3}<\frac{6u}{d^{2}}$, and since $u+1\leq\ell_{1}+\ell_{2}+\ell_{3}$, we have that
$\ell_{1}d+(1+o(1))sd+e(B,B^{C})+wd-w\geq du-\frac{6u}{d}+|B^{C}|\ln^{2}d\geq du+\frac{u\ln^{2}d}{2}.$
We have at most $\binom{n}{u}$ ways to choose the obstruction. Thus, the probability of having an obstruction with such $u,\ell_{1},w,s,b$ is at most
$\binom{n}{u}(1-p)^{u\ln^{2}d/2}$ $\leq\left(\frac{en}{u}\right)^{u}\exp\left\{-\frac{pu\ln^{2}d}{2}\right\}$
$\leq\exp\left\{u\left(1+\ln n-\ln u-\frac{p\ln^{2}d}{2}\right)\right\}$
$\leq\exp\left\{u\left(1+\frac{C^{3}}{p}\ln d-\frac{p\ln^{2}d}{2}\right)\right\}\leq\exp\left\{-\frac{pu\ln^{2}d}{3}\right\},$
where in the penultimate inequality we used $u\geq\frac{n}{d^{C^{3}/p}}$. Union bound over the at most $n^{5}$ choices of $u,\ell_{1},w,s$ and $b$ completes the proof. ∎
@@
Whp there are no obstructions with $b>\frac{n}{2}$ and $e(B^{C},B)\leq|B^{C}|\ln^{2}d$.

### v3

@@ 3.3 Proof of Theorem 3b
where we used that $u\geq d$. Union bound over the less than $n^{2}$ choices of values for $s,w$ completes the proof. ∎
Obstructions with $u\geq\frac{n}{d^{C^{5}/p}}$.
Having treated all obstructions with an upper bound on $u$ we turn to arbitrarily large obstructions. Here we will focus on the set $B$ and distinguish whether $b\leq\frac{n}{2}$ or $b>\frac{n}{2}$.
Before looking at these cases let us remark that again the probability of having an obstruction with given properties is maximised when $s=0$ and $w=0$. Indeed, by Lemma 3.4 the number of obstructions given $u,w,s,\ell_{2},b$ and $\ell_{3}$ is bounded by $2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$. Consider again
$\hat{e}$ $=\ell_{1}d+\sum_{i=1}^{w/2}e(W_{i},W_{i}^{C})+\sum_{i=1}^{\ell_{2}}e(S_{i},S_{i}^{C})+\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})$
$\geq d\ell_{1}+w(d-1)+f^{*}_{S}+\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C}).$
Then, the probability of having such an obstruction is bounded by
$2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}(1-p)^{(\hat{e}-du)/2}.$ (7)
We claim that (7) is decreasing as both $w$ and $s$ are increasing and thus we assume $w=s=0$ in the remaining part of the section (and union bound over the possible values for $w,s$ in the end).
To see this, observe that for each component of size two in $W$, we have
$n(ed)^{2}(1-p)^{2(d-1)}=\frac{(\omega ed)^{2}}{n}(1-p)^{-2}=o(1).$
Similarly, from 3.7 we know that the edge-boundary of $S$ is minimised when all but one component are of size three. For components with three vertices the calculation above holds verbatim, i.e.,
$n(ed)^{3}(1-p)^{3d-3}=\frac{(\omega ed)^{3}}{n^{2}}(1-p)^{-3}=o(1).$
The remaining component of $S$ is of size at most $\frac{n}{d^{C^{3}/p}}$ (by definition of $S$) and for such a component we obtain
$f^{*}\left(\frac{n}{d^{C^{3}/p}}\right)\geq\frac{eC^{2}}{p}\ln d\frac{n}{d^{C^{3}/p}},$
where we used 1.3. Thus, the remaining component of size $s_{max}$ in $S$ expands by a factor of at least $\frac{eC^{2}}{p}\ln d$ and thus we obtain
$n(ed)^{s_{max}}(1-p)^{(eC^{2}/p)\ln d\cdot s_{max}}=\exp(\ln n+s_{max}\ln ed-eC^{2}s_{max}\ln d)=o(1).$
For the remainder of this section fix $u,w,s,\ell_{2},b$, and $\ell_{3}$, and let $\mathcal{B}_{u,w,s,b}$ be the event that there is a minimal obstruction with $u=|U|$, $w=|W|$, $s=|S|$ and $0\neq b=|B|$. Recall that we will bound the probability for $w=s=0$ and take a union bound over the possible values for $w,s$ in the end. Let us start with $b\leq\frac{n}{2}$ and first treat this case under the assumption that the edge boundary of $B$ in $G$ is not too small.
@@
Whp, there are no obstructions with $0\neq b\leq n/2$ and $e(B,B^{C})\geq b\ln^{2}d$.
@@
Similarly to before, we have that
$\hat{e}\geq\ell_{1}d+e(B,B^{C})\geq\ell_{1}d+b\ln^{2}d.$
Now, by Lemma 3.4, there are at most
$2n^{w/2+2\ell_{2}+2\ell_{3}}(ed)^{w/2+s+b}$
such minimal obstructions. Then, we obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2n^{2\ell_{3}}(ed)^{b}(1-p)^{(\hat{e}-du)/2}$
$\leq 2n^{2\ell_{3}}(ed)^{b}(1-p)^{(\ell_{1}d+n\ln^{2}d-du)/2},$
Using that $(1-p)^{d}=\omega/n$, we obtain
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2n^{2\ell_{3}}(ed)^{s+b}\left(\frac{\omega}{n}\right)^{1/2(\ell_{1}-u)}\exp\left(-\frac{p}{2}b\ln^{2}d\right)$
$\leq 2n^{3/2\ell_{3}+1/2(u-\ell)}\omega^{1/2(\ell_{1}-u)}\exp\left((s+b)\ln(ed)-\frac{p}{2}b\ln^{2}d\right)$
$\leq 2\exp\left(\frac{1}{2}(u-\ell)\ln n+\frac{3}{2}\ell_{3}\ln n+\frac{1}{2}(\ell_{1}-u)\ln\omega+b\ln ed-\frac{p}{2}b\ln^{2}d\right)$
where we used $\ell_{1}+\ell_{3}=\ell$. Since $U$ is an obstruction we have $\ell>u$. Furthermore, note that $\ell_{3}\ll\frac{b}{d^{2}}$ such that $\ell_{3}\ln n\ll\frac{b}{d}$ and recall that $\omega$ tends to infinity arbitrarily slowly. Thus, we obtain
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq\exp\left(b\ln ed-\frac{p}{2}b\ln^{2}d\right)$
$\leq\exp\left(-\Omega\left(d^{2}\ln^{2}d\right)\right),$
@@
where we used that $b\geq\frac{n}{d^{C^{3}/p}}\gg d^{2}$. By the union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$, we have that the probability of an obstruction violating the statement of the claim is at most
@@
Whp, there are no obstructions with $0\neq b\leq\frac{n}{2}$ and $e(B,B^{C})\leq b\ln^{2}d$.
@@
Note that every set $B\subseteq V$ with $e(B,B^{C})\leq b\ln^{2}d$ is of size at least $\frac{n}{d^{\ln^{2}d}}$ since by 1.3, for $b<\frac{n}{d^{\ln^{2}d}}$,
$e(B,B^{C})\geq\frac{b\ln^{2}d\log_{C}d}{C}>b\ln^{2}d.$
We again have that
$\hat{e}\geq\ell_{1}d+f^{*}(b).$
@@
Furthermore, since we assume $e(B,B^{C})\leq b\ln^{2}d$ and $b\geq\frac{n}{d^{\ln^{2}d}}$, by Lemma 3.5 there are at most $\exp\left\{\frac{2b}{\ln d}\right\}$ choices for $B$. Therefore, by Lemma 3.3, there are at most $2\exp\left\{\frac{2b}{\ln d}\right\}$ such minimal obstructions. We obtain that
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left(\frac{2b}{\ln d}\right)(1-p)^{(\hat{e}-du)/2}$
$\leq 2\exp\left\{\frac{2b}{\ln d}\right\}(1-p)^{\frac{1}{2}d(\ell_{1}-u)+\frac{f^{*}(b)}{2}}$
$\leq 2\exp\left\{\frac{2b}{\ln d}-\frac{pf^{*}(b)}{2}\right\}\left(\frac{\omega}{n}\right)^{\frac{1}{2}(\ell_{1}-u)}$
$\leq 2\exp\left\{-\frac{1}{2}(\ell_{1}-u)\ln n+\frac{2b}{\ln d}-\frac{pf^{*}(b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}.$
Recall that $\ell=\ell_{1}+\ell_{3}$. Thus,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{-\frac{1}{2}(\ell-u)\ln n+\frac{1}{2}\ell_{3}\ln n+\frac{2b}{\ln d}-\frac{pf^{*}(b)}{2}+\frac{1}{2}(\ell_{1}-u)\ln\omega\right\}$
$\leq 2\exp\left\{\frac{3b}{\ln d}-\frac{pf^{*}(b)}{2}\right\}$
@@
where in the last step we used that $u<\ell$, $\ell_{3}\ll\frac{b}{d^{2}}$ and that $\omega$ tends to infinity arbitrarily slowly. Since we assume that $b\leq\frac{n}{2}$ and $n\leq C^{d}$, we have that $\log_{C}b\leq\log_{C}(C^{d}/2)\leq d-\frac{1}{C}$. Hence, $f^{*}(b)\geq\frac{b}{C^{2}}$. Therefore,
$\mathbb{P}\left(\mathcal{B}_{u,w=0,s=0,b}\right)$ $\leq 2\exp\left\{b\left(\frac{3}{\ln d}-\frac{p}{2C^{2}}\right)\right\}\leq\exp\left\{-\frac{d^{2}}{3C^{2}}\right\}.$
union bound over the at most $n^{6}$ values for $u,w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
Now we turn to obstructions with $b>\frac{n}{2}$. The argument is similar to the other case with a focus on $B^{C}$ and again we split into cases depending on the size of the edge-boundary of $B^{C}$. Since in this regime Lemma 3.9 does not hold, we require a more general (and very crude) bound on $\ell_{3}$.
Claim 3.18.
Whp, there are no obstructions such that $\ell_{3}\geq\frac{9u}{d^{2}}\text{ and }u\geq d^{2}.$
Proof.
Suppose that $u\geq d^{2}$ and $\ell_{3}\geq\frac{9u}{d^{2}}$. Then
$\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})$ $\geq f^{*}_{B}\geq(\ell_{3}-1)\frac{n\ln d}{d^{C^{3}/p}}\geq\left(\frac{9u}{d^{2}}-1\right)\frac{n\ln d}{d^{C^{3}/p}}\geq 8du.$
Recalling that choosing $U$ determines the obstruction, the probability of an obstruction with $\sum_{i=1}^{\ell_{3}}e(B_{i},B_{i}^{C})\geq 8du$ is at most
$\binom{n}{u}(1-p)^{1/2(8du-du)}\leq n^{u}\left(\frac{\omega}{n}\right)^{3.5u}\leq n^{-2.4u},$
where we used $(1-p)^{d}=\omega/n$ for $\omega$ tending to infinity arbitrarily slowly. A union bound over all possible values of $u$, and the at most $n^{2}$ choices for $b$ and $\ell_{3}$ yields that the probability of an event violating the statement of the claim is at most
$3n^{2}\sum_{u=1}^{n}n^{-2.4u}\leq 3n^{2}\cdot 2n^{-2.4}=o(1).$
@@
Note that by Lemma 3.15, we may assume that $u\geq\frac{n}{d^{C^{5}/p}}$, and thus $|B^{C}|\geq\frac{n}{d^{C^{5}/p}}$. We begin by assuming that $B^{C}$ has a large edge-boundary.
@@
Whp, there are no obstructions with $0\neq b>\frac{n}{2}$ and $e(B^{C},B)\geq|B^{C}|\ln^{2}d$.
@@
Again, we consider
$\hat{e}\geq\ell_{1}d+e(B,B^{C})\geq\ell_{1}d+|B^{C}|\ln^{2}d.$
By Claim 3.18, we may assume that $\ell_{3}<\frac{9u}{d^{2}}$, and since $u+1\leq\ell_{1}+\ell_{3}$, we have that
$\hat{e}\geq du-\frac{9u}{d}+|B^{C}|\ln^{2}d\geq du+\frac{u\ln^{2}d}{2}.$
As there are at most $du$ edges touching $U$ in $G$, there are at least $\frac{u\ln^{2}d}{4}$ edges which are not present in $G_{p}$ since we are potentially double-counting any edge between different components outside of $U$. We have at most $\binom{n}{u}$ ways to choose the obstruction. Thus, the probability of having such an obstruction is at most
$\binom{n}{u}(1-p)^{u\ln^{2}d/4}$ $\leq\left(\frac{en}{u}\right)^{u}\exp\left\{-\frac{pu\ln^{2}d}{4}\right\}$
$\leq\exp\left\{u\left(1+\ln n-\ln u-\frac{p\ln^{2}d}{4}\right)\right\}$
$\leq\exp\left\{u\left(1+\frac{C^{5}}{p}\ln d-\frac{p\ln^{2}d}{4}\right)\right\}\leq\exp\left\{-\frac{pu\ln^{2}d}{5}\right\},$
where in the penultimate inequality we used $u\geq\frac{n}{d^{C^{5}/p}}$. Union bound over the at most $n^{6}$ choices of $u,\ell_{1},w,s,\ell_{2},b$ and $\ell_{3}$ completes the proof. ∎
@@
Whp, there are no obstructions with $0\neq b>\frac{n}{2}$ and $e(B^{C},B)\leq|B^{C}|\ln^{2}d$.